# WaterSplatting: Fast Underwater 3D Scene Reconstruction Using Gaussian Splatting

### v1

@@ 2.3 Computer Vision in Scattering Media
There are many challanges in underwater computer vision. The complex lighting conditions including scattering and attenuation of light leading to distorted images and the failure of traditional algorithms trained on clear-air scenes [12]. WaterNeRF [33] estimates the medium parameters separately from the rendering and works with images that have been histogram-equalized. In our method we directly modify the rendering equation. ScatterNeRF [30] uses a volumetric representation of the scene, by extending the NeRF rendering equation to include scattering properties of the medium for adverse weather environments, with a separation of the backscatter component of the image. Our method is more specialized for underwater and foggy scenes. SeaThru [2] introduces a method for removing water from underwater images. This method addresses color distortion in underwater images by revising the image formation model from [1], accurately estimating backscatter, and correcting colors along the depth axis. SeaThru-NeRF [18] implements the image formation model [1] that separates direct and backscatter components, into the NeRF rendering equations, which is highly specialized for underwater scenes. We implemented a similar model but on Gaussian Splatting, which yields higher performance and enables real-time rendering.
@@ 3.1 Preliminaries
3DGS primitive also has two additional parameterized properties: opacity and spherical harmonics coefficients to represent directional appearance component (anisotropic color). In order to render pixel-wise color, primitives are transformed into camera space via a viewing transformation and the Jacobian of the affine approximation of the projective transformation on , then we get the projected 2D means and 2D covariance matrices as:
@@
where is the coordinate of the pixel. For rasterization, each Gaussian is truncated at 3 sigma, considering only those intersecting with the patch comprising pixels within this range, based on the property that about of the probability lies within 3 sigma of the mean. Thus, the pixel color is computed by alpha blending of the sorted intersected Gaussians whose are higher than a threshold:
@@ 3.3 Loss Function Alignment
In vanilla 3DGS, the loss function is combined with an Loss and a D-SSIM Loss, which suits primitives without shared parameters. Inspired by [25], we propose a regularized loss function to boost the weight of the dark regions in optimization to align with how humans perceive dynamic range. To be more specific, we apply a pixel-wise weight on both rendered estimate and target image , where with pixel coordinate and denotes stopping gradient of its argument, which backpropagates zero derivative.
Thus, we have our regularized Loss
which results in less blur and sharper edges, and our regularized D-SSIM Loss
which encourages high structural similarity between and . Our final proposed loss function is
which encourages Gaussian optimization to better align the human eye’s perception of the dynamic range.
@@ 4.1 Experiments
| SeaThru-NeRF | 25.908 | 0.785 | 0.304 | 30.193 | 0.873 | 0.210 | 21.841 | 0.767 | 0.249 | 27.846 | 0.834 | 0.224 | 0.07 | 10h |
| SeaThru-NeRF-NS | 26.755 | 0.826 | 0.168 | 30.959 | 0.915 | 0.133 | 23.282 | 0.876 | 0.111 | 31.276 | 0.937 | 0.071 | 0.9 | 2h |
| ZipNeRF | 16.937 | 0.474 | 0.412 | 19.956 | 0.442 | 0.421 | 19.022 | 0.349 | 0.483 | 19.012 | 0.349 | 0.482 | 0.9 | 6h |
@@
| Ours | 29.840 | 0.889 | 0.203 | 32.203 | 0.948 | 0.116 | 24.741 | 0.892 | 0.116 | 31.616 | 0.942 | 0.080 | 41.8 | 9.4min |
Implementation Details: Our implementation is based on the reimplemented version of 3DGS released by NeRF-Studio. Following [42, 46], we accumulate the norms of the individual pixel gradients of for primitive densification.
For the medium encoding, we use a spherical harmonic encoding [37] and a MLP with 2 linear layers with 128 hidden units and Sigmoid activation, followed by Sigmoid activation for and Softplus activation for and . At each densification and pruning step of the 3DGS, the moving averages in the Adam optimizer of the medium encoding are reset.
SeaThru-NeRF Dataset: SeaThru-NeRF Dataset released by [18] contains real-world scenes acquired from four different scenes in sea: IUI3 Red Sea, Curaçao, Japanese Gardens Red Sea, and Panama. There are 29, 20, 20 and 18 images respectively, among which 25, 17, 17 and 15 images are used for training and the rest 4, 3, 3 and 3 are used for validation. The dataset encompasses a variety of water and imaging conditions. The images were captured in RAW format using a Nikon D850 SLR camera housed in a Nauticam underwater casing with a dome port, which helped prevent refractions that could disrupt the pinhole model. These images were then downsampled to an approximate resolution of 900 × 1400. Prior to processing, the input linear images underwent white balancing with a clipping per channel to eliminate extreme noise pixels. Lastly, COLMAP [32] was employed to determine the camera poses.
@@ 4.2 Results
First, we evaluated the performance of our method using the standard benchmark dataset, the SeaThru-NeRF Dataset. Table 1 compares PSNR, SSIM, LPIPS, average FPS and average training time with the baseline methods across the validation sets of four scenes. Our method demonstrates its superiority in the majority of cases and efficiency on both rendering and training.
@@
Fig. 4 demonstrates that in both the ’IUI3 Red Sea’, the ’Japanese Gardens Red Sea’ and ’Panama’ scenes, our method delivers superior quality and more effectively separates medium and object than the SeaThru-NeRF, especially in deeper and more complex scenes (as highlighted in the red square). Additionally, our depth map reveals much finer details compared to SeaThru-NeRF, which struggles to produce a reasonable depth map at greater distances, as indicated by the red color in the upper right corner of the depth map. We also achieve higher PSNR values in both scenes. The same advantages are observed in simulated scenes, where our method renders better details (indicated by the red square) than the SeaThru-NeRF in both easy and hard foggy scenes as depicted in Fig. 5. Our rendering without medium and depth maps significantly outperform those from the SeaThru-NeRF, especially in scenes that are farther from the camera. While our method’s predictions may appear blurry and the object map unclear in the upper right corner for the hard foggy scene, the results from SeaThru-NeRF are considerably worse. The restoration quality comparison presented in Table 2 further quantitatively demonstrates the superiority of our method in simulated scenes. Overall, our method surpasses SeaThru-NeRF in both underwater and simulated scenes.
@@ 4.3 Ablation Study
We isolate the different contributions to address the key one leading to the increment. We conduct a quantitative analysis on different combination of loss functions, between pixel-wise component {, , } and frame-wise {, }, as well as removing the medium effect and removing both the medium and our proposed loss function as a reimplementation of 3DGS. These comparisons are made across validation sets for the SeaThru-NeRF dataset in Table 3.
Fig. 6 clearly shows that our frame-level provides better details in low light. When used alone, can also improve the reconstruction quality and structural similarity of distant objects. Our pixel-level can further sharpen the edges. The supervision ability of [25] for 3DGS-based models is relatively insufficient. Our proposed shows superiority over other configurations in leading the 3DGS-based model to better fit HDR scenes and removing the medium component (basically 3DGS) significantly hurms the performance of our method, which indicates the necessity of our approach.

### v2

@@ 2.3 Computer Vision in Scattering Media
There are many challanges in underwater computer vision. The complex lighting conditions including scattering and attenuation of light leading to distorted images and the failure of traditional algorithms trained on clear-air scenes [12]. SeaThru [2] introduces a method for removing water from underwater images. This method addresses color distortion in underwater images by revising the image formation model from [1], accurately estimating backscatter, and correcting colors along the depth axis. WaterNeRF[33] estimates medium parameters separately from rendering, while ScatterNeRF[30] extends NeRF’s volumetric rendering to model scattering in adverse weather. [zhang2023beyond] introduces a neural reflectance field to jointly learn scene albedo, normals, and medium effects, achieving color consistency by modeling light attenuation and backscatter through logistic regression and differentiable volume rendering. [tang2024uwnerf] explores hybrid neural-explicit approaches for reconstructing dynamic underwater environments, though its reliance on volumetric rendering limits real-time applicability. SeaThru-NeRF [18] implements the image formation model [1] that separates direct and backscatter components, into the NeRF rendering equations, which is highly specialized for underwater scenes. We implemented a similar model but on Gaussian Splatting, which yields higher performance and enables real-time rendering.
@@ 3.1 Preliminaries
3DGS primitive also has two additional parameterized properties: opacity and spherical harmonics coefficients to represent directional appearance component (anisotropic color). In order to render pixel-wise color, primitives are transformed into camera space via a viewing transformation , the Jacobian of the affine approximation of the projective transformation on and a projection matrix . Then we get the projected 2D means and 2D covariance matrices :
@@
where is the coordinate of the pixel. For rasterization, each Gaussian is truncated at 3 sigma, considering only those intersecting with the patch comprising pixels within this range, based on the property that about of the probability lies within 3 sigma of the mean. Pixel colors are computed by alpha blending of the sorted intersected Gaussians whose are higher than a threshold:
@@ 3.3 Loss Function Alignment
In vanilla 3DGS, the loss function is a combination of loss and D-SSIM loss. In low-light situations, [25] proposed a regularized loss
to boost the weight of the dark regions in optimization to align with how humans perceive dynamic range, which is applied to underwater scene recontruction by [18]. For the case of our 3DGS-based model, we propose a regularized loss function : we apply pixel-wise weight on both rendered estimate and target image , where with pixel coordinate and denotes stopping gradient of its argument, which backpropagates zero derivative.
By introducing the weight to the loss function of 3DGS composing of and D-SSIM loss, we have the regularized loss
and the regularized D-SSIM loss
prioritizes structural similarity and perceptual fidelity, particularly in dark regions where human perception is more sensitive. Integrating regularization into the formulation becomes particularly critical for 3DGS optimization due to the discrete nature of its primitives, necessitating structural regularization to maintain perceptual consistency across independently optimized elements - a distinct requirement from NeRF-based approaches where parameter-shared volumetric representations can achieve adequate low-light adaptation through pixel-level regularized losses alone. To model the smoothness of volumetric medium, we employ as our pixel-level loss. Our final proposed loss function is
@@ 4 Experiments
SeaThru-NeRF | 27.310 | 0.861 | 0.221 | 29.305 | 0.898 | 0.214 | 22.116 | 0.802 | 0.232 | 28.608 | 0.868 | 0.202 | 0.07 | 10h |
SeaThru-NeRF-NS | 27.651 | 0.885 | 0.133 | 30.609 | 0.928 | 0.159 | 23.542 | 0.889 | 0.116 | 31.857 | 0.948 | 0.071 | 0.9 | 2h |
ZipNeRF | 29.350 | 0.899 | 0.106 | 29.925 | 0.938 | 0.124 | 23.446 | 0.883 | 0.136 | 32.335 | 0.956 | 0.064 | 0.9 | 6h |
@@
Ours | 29.386 | 0.910 | 0.180 | 32.672 | 0.957 | 0.110 | 25.201 | 0.904 | 0.113 | 31.490 | 0.948 | 0.075 | 41.8 | 9.4min |
Implementation Details: Our implementation is based on the reimplemented version of 3DGS released by NeRF-Studio [ye2024gsplat]. Following [43, 47], we accumulate the norms of the individual pixel gradients of for primitive densification.
For the medium encoding, we use a spherical harmonic encoding [38] and a MLP with 2 linear layers with 128 hidden units and Sigmoid activation, followed by Sigmoid activation for and Softplus activation for and . Upon each densification and pruning step of the 3DGS, the moving averages in the Adam optimizer of the medium encoding are reset to ensure the independence of subsequent iterations. We reset opacity to 0.5 every 500 training steps and prune gaussians with opacity below 0.5 every 100 training steps.
The culling and reset thresholds [15] are set to 0.5 to encourage Gaussians to model opaque objects.
SeaThru-NeRF Dataset: SeaThru-NeRF Dataset released by [18] contains real-world scenes acquired from four different scenes in sea: IUI3 Red Sea, Curaçao, Japanese Gardens Red Sea, and Panama. There are 29, 20, 20 and 18 images respectively, among which 25, 17, 17 and 15 images are used for training and the rest 4, 3, 3 and 3 are used for validation. The dataset encompasses a variety of water and imaging conditions. The images were captured in RAW format using a Nikon D850 SLR camera housed in a Nauticam underwater casing with a dome port, which helped prevent refractions that could disrupt the pinhole model. These images were then downsampled to an approximate resolution of 900 × 1400. Prior to processing, the input linear images underwent white balancing with a clipping per channel to eliminate extreme noise pixels. Lastly, COLMAP [32] was employed to determine the camera poses and correct lens distortions inherent.
@@ 4.1 Results
First, we evaluated the performance of our method using the standard benchmark dataset, the SeaThru-NeRF Dataset. Table 1 compares PSNR, SSIM, LPIPS, average FPS and average training time with the baseline methods across the validation sets of four scenes. Our method demonstrates its superiority in the majority of cases and efficiency on both rendering and training. ’Panama’ is a special case where there is little medium and the medium properties stays the same across different ray directions. Therefore, ZipNeRF can reconstruct it well. However, ZipNeRF training takes orders of magnitude more time than our method and does not offer real-time rendering.
@@
Fig. 4 demonstrates that in the ’IUI3 Red Sea’, the ’Japanese Gardens Red Sea’ and ’Panama’ scenes, our method delivers superior quality and more effectively separates medium and object than the SeaThru-NeRF, especially in deeper and more complex scenes (as highlighted in the red square). Additionally, our depth map reveals much finer details compared to SeaThru-NeRF, which struggles to produce a reasonable depth map at greater distances, as indicated by the red color in the upper right corner of the depth map. We also achieve higher PSNR values in both scenes. The same advantages are observed in simulated scenes, where our method renders better details (indicated by the red square) than the SeaThru-NeRF in both easy and hard foggy scenes as depicted in Fig. 5. Our rendering without medium and depth maps significantly outperform those from the SeaThru-NeRF, especially in scenes that are farther from the camera. While our method’s predictions may appear blurry and the object map unclear in the upper right corner for the hard foggy scene, the results from SeaThru-NeRF are considerably worse. The restoration quality comparison presented in Table 2 further quantitatively demonstrates the superiority of our method in simulated scenes. Overall, our method surpasses SeaThru-NeRF in both underwater and simulated scenes.
@@ 4.2 Ablation Study
We isolate the different contributions to show their importance. We conduct a quantitative analysis on different combination of loss functions, between pixel-wise component {, , , } and frame-wise {, }, as well as removing the medium effect and removing both the medium and our proposed loss function as a reimplementation of 3DGS. These comparisons are made across validation sets for the SeaThru-NeRF dataset in Table 3.
Fig. 6 clearly demonstrates that our frame-level improves the reconstruction quality and structural similarity of distant objects in low light. When used alone, can also provide better details in far distance. Used with , pixel-level and can further enhance reconstruction quality. However, alone [25] is not sufficient to train 3DGS-based models, and is required. Our proposed shows superiority over other configurations in leading the 3DGS-based model to better fit HDR scenes and removing the medium component (basically 3DGS) significantly hurms the performance of our method, which indicates the necessity of our approach.